solve.
s^2+3s-54=0

Answer by ewatrrr(24785)


Hi,
solving
s^2+3s-54=0
factoring
(s+9)(s-6) = 0 Note:SUM of the inner product(9s) and the outer product(-6s) = 3s
(s+9)(s-6)= 0
s = -9
(s-6) = 0
s = 6